# Improved High-probability Convergence Guarantees of Decentralized SGD

### v2

@@ Appendix I Achieving Linear Speed-up
Linear speed-up for non-convex costs. The crucial result for obtaining linear speed for non-convex costs is Lemma 2, which establishes the variance reduction benefits of decentralized methods in the HP sense. To see why this is the case, note that using a fixed step-size $\alpha_{t}\equiv\alpha$ and following the same steps as in the proof of Theorem 1, while relying only on assumption (A4) to control the noise, one would get a final bound where the leading term is of the form
@@
where it is easy to see that choosing $\alpha=\Theta\Big(\frac{\sqrt{n}}{\sqrt{T}}\Big)$ leads to the desired linear speed-up. Works like [dsgd-high-prob, dsgd-online-noncooperative-games, dsgd-online-stochastic] provide no results on the variance reduction benefit of decentralized learning in the HP sense, instead relying only on the definition of sub-Gaussianity. As such, even if a fixed step-size was used, the authors would be unable to obtain a bound of the form (50) and achieve linear speed-up. As discussed in Appendix E, using a time-varying step-size for non-convex costs inevitably leads to a loss of linear speed-up, as all terms decay at the same rate, leading to a bound of the form
@@
where the term not achieving linear speed-up is of higher order. This is consistent with MSE guarantees, e.g., [pmlr-v119-koloskova20a, ran-improved, unified-refined], where a fixed step-size is also needed to achieve linear speed-up.
Linear speed-up for strongly convex costs. For strongly convex costs, in addition to Lemma 2, a tighter control of the MGF of the process of interest is required, which is achieved in Lemma 5. To see why this is the case, consider the existing techniques in centralized settings, e.g., [pmlr-v99-harvey19a, pmlr-v206-bajovic23a], where the authors show via induction that the MGF of the process of interest is uniformly bounded, i.e., that there exists a $B>0$, such that, for all $t\geq 1$
$\mathbb{E}[\exp(BF^{t})]\leq\exp(1).$ (51)
@@
$\mathbb{P}\bigg(f(\overline{x}^{t})-f^{\star}>\frac{\epsilon+1}{nB(t+t_{0})}\bigg)=\mathbb{P}\big(BF^{t}>\epsilon+1\big)\leq\exp(-\epsilon),$
or equivalently, with probability at least $1-\delta$
@@
$\mathbb{E}[\exp(BF^{1})]=\exp\Big(nB(t_{0}+1)\big(f(\overline{x}^{1})-f^{\star}\big)\Big)\leq\exp(1),$
where the equality follows from the definition of $F^{t}$ and the fact that the initial models are chosen deterministically, while the inequality follows if and only if $B\leq\frac{1}{n(t_{0}+1)\big(f(\overline{x}^{1})-f^{\star}\big)}$. As such, in order for (51) to hold, the constant $B$ has to decay with $n$, canceling the linear speed-up in (52) and resulting in a HP bound of the form
@@
for some problem related constants $G_{i}>0$, $i\in[4]$, providing a much tighter and more fine-grained control of the MGF compared to (52) and crucially allowing for a bound where the leading term achieves linear speed-up in the number of users (as $G_{1}$ is independent of $n$). On the other hand, using a different approach to [pmlr-v99-harvey19a, pmlr-v206-bajovic23a], the work [liu2024revisiting] establishes the following HP bound for the last iterate of SGD in centralized setting and smooth, strongly convex costs
@@
While such result could potentially be leveraged in decentralized settings to maintain linear speed-up (provided that the constant growing linearly with $n$ affects only the higher-order term), it still accrues a factor of $\log(t)$ in the leading term, whereas in our bound it appears only in the higher-order term, decaying at rate $\mathcal{O}\Big(\frac{\log(t+t_{0})}{(t+t_{0})^{3}}\Big)$. As such, Lemma 5 provides the most fine-grained control of the MGF and is of independent interest when studying “almost decreasing” type processes, even in centralized settings.
@@
As discussed in the main body, to model the communication network, we use assumption (A1), which subsumes fixed undirected networks, as well as a class of fixed directed networks. The authors in [dsgd-high-prob] consider a more general case of directed, time-varying networks, facilitated by the following condition.
@@
Our work can be readily extended to the setting of time-varying networks, with assumption (A1) replaced by (A1). To see why this is the case, note that in our proofs we only use the fact that under (A1), we have $\lambda=\|W-J\|\in[0,1)$. The authors in [dsgd-high-prob, Lemma 1] use a similar result for time-varying networks, namely that under (A1), for any time indices $t\geq s$
@@
where $\Phi(t,s)=W^{t-1}\times\ldots\times W^{s}$ is the transition matrix from time $s$ to $t$ (with $\Phi(t,t)=I_{n}$), while $M>0$ and $\rho\in(0,1)$ are graph related constants. As discussed in [dsgd-high-prob], the constant $\rho$ plays the same role of the network connectivity parameter, as $\lambda$ does in our work. It can be readily verified that replacing (A1) with (A1) and using (53) to control the consensus gap, our results readily go through, with $\lambda$ being replaced by $\rho$. As such, our results can be easily extended to the (directed) time-varying network setting.
@@
see [jin2019short, Lemma 1]. This argument allows us to show the variance reduction benefit of decentralized learning, i.e., that $\overline{z}^{t}$ is $\mathcal{O}\Big(\frac{\sigma\sqrt{d}}{\sqrt{n}}\Big)$-sub-Gaussian (in the sense of (A4)), at the cost of introducing a $\sqrt{d}$ dependence. One way to mitigate this is to directly assume that the noise $z^{t}_{i}$ at each user is zero-mean and $\sigma_{i}$-norm-sub-Gaussian. Using a similar argument to the one in Lemma 6 and Corollary 7 in [jin2019short], we could then show that $\overline{z}^{t}$ is $\frac{\sigma}{n}$-sub-Gaussian, while reducing the dependence on problem dimension to $\log(d)$. This, however, comes at the cost of imposing a slightly stronger noise condition, as any $\sigma$-norm-sub-Gaussian random vector is also $\sigma$-sub-Gaussian (in the inner product sense), while the opposite implication inevitably introduces a $\sqrt{d}$ factor, as is the case in Lemma 2. Finally, we note that, in the regime $n\ll d$, we can simply use the above argument following Sedrakyan’s inequality, to conclude that $\overline{z}^{t}$ is $\sigma$-sub Gaussian, reducing the dependence on problem dimension $d$, at the cost of losing linear speed-up in the number of users.

### v3

@@ Appendix I Achieving Linear Speed-up
Linear speed-up for non-convex costs.
The crucial result for obtaining linear speed for non-convex costs is Lemma 2, which establishes the variance reduction benefits of decentralized methods in the HP sense. To see why this is the case, note that using a fixed step-size $\alpha_{t}\equiv\alpha$ and following the same steps as in the proof of Theorem 1, while relying only on assumption (A4) to control the noise, one would get a final bound where the leading term is of the form
@@
where it is easy to see that choosing $\alpha=\Theta\Big(\frac{\sqrt{n}}{\sqrt{T}}\Big)$ leads to the desired linear speed-up. Works like [44, 45, 76] provide no results on the variance reduction benefit of decentralized learning in the HP sense, instead relying only on the definition of sub-Gaussianity. As such, even if a fixed step-size was used, the authors would be unable to obtain a bound of the form (52) and achieve linear speed-up. As discussed in Appendix E, using a time-varying step-size for non-convex costs inevitably leads to a loss of linear speed-up, as all terms decay at the same rate, leading to a bound of the form
@@
where the term not achieving linear speed-up is of higher order. This is consistent with MSE guarantees, e.g., [33, 75, 1], where a fixed step-size is also needed to achieve linear speed-up.
Linear speed-up for strongly convex costs.
For strongly convex costs, in addition to Lemma 2, a tighter control of the MGF of the process of interest is required, which is achieved in Lemma 5. To see why this is the case, consider the existing techniques in centralized settings, e.g., [21, 5], where the authors show via induction that the MGF of the process of interest is uniformly bounded, i.e., that there exists a $B>0$, such that, for all $t\geq 1$ and $\nu\in(0,B]$
$\mathbb{E}[\exp(\nu F^{t})]\leq\exp\Big(\frac{\nu}{B}\Big).$ (53)
@@
$\mathbb{P}\bigg(f(\overline{x}^{t})-f^{\star}>\frac{\epsilon+1}{n\nu(t+t_{0})}\bigg)=\mathbb{P}\big(\nu F^{t}>\epsilon+1\big)\leq\exp(-\epsilon).$
Choosing $\nu=B$, it follows that, for any $\delta\in(0,1)$, with probability at least $1-\delta$
@@
$\mathbb{E}[\exp(\nu F^{1})]=\exp\Big(\nu n(t_{0}+1)\big(f(\overline{x}^{1})-f^{\star}\big)\Big)\leq\exp\Big(\frac{\nu}{B}\Big),$
where the equality follows from the definition of $F^{t}$ and the fact that the initial models are chosen deterministically, while the inequality follows if and only if $B\leq\frac{1}{n(t_{0}+1)(f(\overline{x}^{1})-f^{\star})}$. As such, in order for (53) to hold, the constant $B$ has to decay with $n$, canceling the linear speed-up in (54) and resulting in a HP bound of the form
@@
for some problem related constants $G_{i}>0$, $i\in[4]$, providing a much tighter and more fine-grained control of the MGF compared to (53) and crucially allowing for a bound where the leading term achieves linear speed-up in the number of users (as $G_{1}$ is independent of $n$). On the other hand, using a different approach to [21, 5], the work [41] establishes the following HP bound for the last iterate of SGD in centralized setting and smooth, strongly convex costs
@@
While such result could potentially be leveraged in decentralized settings to maintain linear speed-up (provided that the constant growing linearly with $n$ affects only the higher-order term), it still accrues a factor of $\log(t)$ in the leading term, whereas in our bound the $\log$ term only affects a higher-order term, decaying at rate $\mathcal{O}\Big(\frac{\log(t+t_{0})}{(t+t_{0})^{3}}\Big)$. As such, Lemma 5 provides the most fine-grained control of the MGF and is of independent interest when studying “almost decreasing” type processes, even in centralized settings.
@@
As discussed in the main body, to model the communication network, we use assumption (A1), which subsumes fixed undirected networks, as well as a class of fixed directed networks. The authors in [44] consider a more general case of directed, time-varying networks, facilitated by the following condition.
@@
Our work can be readily extended to the setting of time-varying networks, with assumption (A1) replaced by (A1). To see why this is the case, note that in our proofs we only use the fact that under (A1), we have $\lambda=\|W-J\|\in[0,1)$. The authors in [44] use a similar result for time-varying in networks (see Lemma 1 [44]), namely that under (A1), for any time indices $t\geq s$
@@
where $\Phi(t,s)=W^{t-1}\times\ldots\times W^{s}$ is the transition matrix from time $s$ to $t$ (with $\Phi(t,t)=I_{n}$), while $M>0$ and $\rho\in(0,1)$ are graph related constants. As discussed in [44], the constant $\rho$ plays the same role of the network connectivity parameter, as $\lambda$ does in our work. It can be readily verified that replacing (A1) with (A1) and using (55) to control the consensus gap, our results readily go through, with $\lambda$ being replaced by $\rho$. As such, our results can be easily extended to the (directed) time-varying network setting.
@@
see Lemma 1 in [27]. This argument allows us to show the variance reduction benefit of decentralized learning, i.e., that $\overline{z}^{t}$ is $\mathcal{O}\Big(\frac{\sigma\sqrt{d}}{\sqrt{n}}\Big)$-sub-Gaussian (in the sense of (A4)), at the cost of introducing a $\sqrt{d}$ dependence. One way to mitigate this is to directly assume that the noise $z^{t}_{i}$ at each user is zero-mean and $\sigma_{i}$-norm-sub-Gaussian. Using a similar argument to the one in Lemma 6 and Corollary 7 in [27], we could then show that $\overline{z}^{t}$ is $\frac{\sigma}{n}$-sub-Gaussian, while reducing the dependence on problem dimension to $\log(d)$. This, however, comes at the cost of imposing a slightly stronger noise condition, as any $\sigma$-norm-sub-Gaussian random vector is also $\sigma$-sub-Gaussian (in the inner product sense), while the opposite implication inevitably introduces a $\sqrt{d}$ factor, as is the case in Lemma 2. Finally, we note that, in the regime $n\ll d$, we can simply use the above argument following Sedrakyan’s inequality, to conclude that $\overline{z}^{t}$ is $\sigma$-sub Gaussian, reducing the dependence on problem dimension $d$, at the cost of losing linear speed-up in the number of users.